A researcher hypothesizes that a person who drops a stack of books and uses profanity will be less likely to gain help (defined as whether or not someone helps pick up the books) than someone who does not use profanity. He flips a coin to randomly assigns participants to each condition, one group says, “Oh darn” and the other says, “Oh s..t” after dropping several books. The researcher then records whether or not a bystander helped pick up the books. The results are listed below. Assume a two-tailed test with .05 level of significance and equal distribution.

bystander helped bystander did not help TOTAL
no profanity actual count 23 16 39
expected count 18.5 20.5 39.0
profanity actual count 14 25 39
expected count 18.5 20.5 39.0
TOTAL actual count 37 41 78
expected count 37.0 41.0 78.0

a) What is the appropriate chi-square test?

b) State the null and alternative hypothesis.

c) In reading the crosstabulation chart, what is the actual amount of no profanity, bystander helped?

d) What is the expected value?
Answer #1

a) Chi square test of independence is appropriate

\chi ^2=\sum \frac{(O_f-E_f)^2}{E_f}=4.16

O is the observed frequency

E is the expected frequency

b) H_0: a person who drops a stack of books and uses profanity will be equally likely to gain help than someone who does not use profanity

H_a:a person who drops a stack of books and uses profanity will be less likely to gain help

c) actual amount is 23

d) Expected value is 18.5